# Expressivity Limits and Trainability Guarantees in Quantum Walk-based Optimization

### v1

@@ Abstract
Quantum algorithms have emerged as a promising tool to solve combinatorial optimization problems. The quantum walk optimization algorithm (QWOA) is one such variational approach that has recently gained attention. In the broader context of variational quantum algorithms (VQAs), understanding the expressivity and trainability of the ansatz has proven critical for evaluating their performance. A key method to study both these aspects involves analyzing the dimension of the dynamic Lie algebra (DLA). In this work, we derive novel upper bounds on the DLA dimension for QWOA applied to arbitrary optimization problems. The consequence of our result is twofold: (a) it allows us to identify complexity-theoretic conditions under which QWOA must be overparameterized to obtain optimal or approximate solutions, and (b) it implies the absence of barren plateaus in the loss landscape of QWOA for $\mathsf{NP}$ optimization problems with polynomially bounded cost functions ($\mathsf{NPO}\text{-}\mathsf{PB}$).
@@ 1 Introduction
Combinatorial optimization problems play an important role in modeling and solving real-world systems. Some common classical approaches to solving these problems include Lagrange multipliers, simulated annealing, branch-and-bound, and evolutionary algorithms [1, 2]. Recent advances in quantum computing have also led to the development of quantum algorithms designed to solve these problems [3, 4, 5]. Among them, variational quantum algorithms (VQAs) [6] — which combine classical optimization with parametrized quantum circuits (PQCs), i.e., quantum circuits composed of unitary operations controlled by tunable parameters — have emerged as a prominent class and are particularly well suited for noisy quantum devices. Arguably, the most prominent VQA for combinatorial optimization is the quantum approximate optimization algorithm (QAOA) [7], a gate-based heuristic inspired by the quantum adiabatic algorithm [8, 9], that alternates between problem-specific and mixing unitaries to explore the solution space. The quantum walk optimization algorithm (QWOA) [10, 11] is a special variant of QAOA where the standard QAOA mixing unitary is replaced with a continuous-time quantum walk (CTQW) [12] operator. QWOA has been proposed as a natural alternative for solving combinatorial optimization problems with structured constraints, where quantum walks can be efficiently implemented through an indexing procedure that maps each feasible solution to a unique computational basis state. Typically, QWOA is implemented using a CTQW on a complete graph [13, 14, 15, 16], which is also the version we consider in this work  [^1] Note that, up to a rescaling of the variational parameter, QWOA is equivalent to a variant of QAOA called the Grover Mixer Quantum Alternating Operator Ansatz (GM-QAOA) [17]. Accordingly, all results and discussions in this work apply directly to that ansatz as well, and vice versa, meaning that results from that ansatz can be directly applied here..
The success of a VQA depends on two key factors. First, the PQC (a.k.a. ansatz) should be expressive enough to prepare the state that encodes the solution to the considered problem. Second, the ansatz should be trainable; that is, it should allow efficient optimization of its parameters to reach such solutions. Several approaches have been proposed to study the expressivity of PQCs, including techniques based on covering numbers [18, 19] and $t$-designs [20]. Moreover, it has been observed that high expressivity in a PQC often comes at the expense of trainability. In particular, it has been shown that if an ansatz forms a $2$-design, it necessarily exhibits barren plateaus [21] — regions in the parameter space where the variance of the loss function and its gradients vanishes exponentially with system size. In such regions, the optimization landscape becomes effectively flat, rendering gradient-based optimization methods inefficient, thus hindering the training of the PQCs.
@@
Yet another measure of expressivity, which we would focus on in this work, is the dimension of the dynamic Lie algebra (DLA) [22, 23, 24, 25, 26]. Given the set of Hamiltonians that generate the unitaries in a PQC, the corresponding DLA is constructed by taking repeated nested commutators of the Hamiltonians. This forms a Lie algebra that characterizes the space of unitaries accessible through time evolution. The dimension of the DLA determines the extent to which the circuit can explore the full unitary group or a restricted subspace, making it a fundamental measure of expressivity. If the DLA spans the full Lie algebra $\mathfrak{su}(d)$ for a $d$-dimensional Hilbert space, the ansatz is considered maximally expressive. However, in many practical scenarios, the DLA forms a lower-dimensional subalgebra, imposing structure on the reachable unitary transformations. As pointed out by Larocca et al. [23], the dimension of the DLA also provides an upper bound for the number of parameters required for a PQC to reach the overparameterization regime, beyond which additional parameters do not enhance the expressivity of the ansatz. Concurrently, Ragone et al. [25] further established that the dimension of the DLA is inversely proportional to the variance of the loss function. This provides a rigorous framework for diagnosing the presence of barren plateaus based on algebraic properties of the circuit.
@@
In the context of QAOA, DLA has been studied both numerically and analytically [22, 23, 24, 26] for the Max-Cut problem, a combinatorial optimization problem defined on graphs and widely studied in the QAOA literature [27]. On the numerical side, the DLA dimension has been evaluated for two specific classes: $2$-regular graphs (also known as cycle graphs or ring of disagrees) and chain graphs (or path graphs), achieving scaling of $\mathcal{O}(n)$ and $\mathcal{O}(n^{2})$, respectively, where $n$ is the number of vertices of the graph [22, 24]. On the analytical side, Allcock et al. [26], using tools from group theory, proved that the exact dimension of the DLA for $2$-regular graphs is indeed linear in $n$, and also established that the exact DLA dimension for the complete graph is in the order of $n^{3}$. These findings, as discussed earlier, have implications both in the context of expressivity and trainability of QAOA ansatz for the Max-Cut problem. For instance, numerical findings by Larocca et al. [23] showed that the number of layers required to solve Max-Cut with a high success probability, on $2$-regular graphs and chains, matches the scaling of their corresponding DLA dimensions  [^2] Note that the number of variational parameters of QAOA/QWOA is linearly related to its number of layers. So, in asymptotic discussions, we can use the number of layers of QAOA/QWOA as a metric for parameterization without loss of generality., thus allowing us to infer that, for these classes of graphs, QAOA does not require overparameterization to achieve optimal solutions.
@@
While DLAs have been extensively studied in the context of QAOA, it remains underexplored in the setting of QWOA. In this work, we initiate a systematic study of DLAs within the QWOA framework, focusing on two key questions: (a) whether overparameterization is necessary for solving or approximating combinatorial optimization problems, and (b) whether barren plateaus arise in the loss landscape.
@@
Another important consequence of our general bound is the absence of barren plateaus in the optimization landscape of QWOA for problems in the class $\mathsf{NPO}\text{-}\mathsf{PB}$ (Theorem 4). We combine our result that QWOA exhibits a polynomially bounded DLA dimension for $\mathsf{NPO}\text{-}\mathsf{PB}$ problems with the aforementioned theoretical framework of Ragone et al. [25], which relates the DLA dimension to the variance of the loss function. This allows us to conclude that, for such problems, the variance of the loss function decays no more than polynomially with input size, thereby ruling out the presence of barren plateaus in this setting. Our finding generalizes the argument of Headley and Wilhelm [28], who showed barren plateau avoidance under asymptotic convergence of the probability distribution associated with the solution space, as our result holds for any $\mathsf{NPO}\text{-}\mathsf{PB}$ problem.
@@
This paper is organized as follows. In Section 2, we introduce some background on combinatorial optimization, complexity classes, QWOA, DLA, and barren plateaus. We present our main results and their theoretical implications in Section 3. Section 4 illustrates these results through the well-known problems of unstructured search, Max-Cut, and $k$-Densest Subgraph. Given the generality of the theoretical results presented, the examples included are not intended to serve as benchmarks, but rather to illustrate broader underlying phenomena. Lastly, we conclude our findings and discuss the future directions of our work in Section 5.
@@ 2.1 Combinatorial Optimization Problems
We consider combinatorial optimization problems defined by a cost function (the objective function) $C:\mathcal{S}\rightarrow\mathbb{R}$, where $\mathcal{S}$ — the solution space — is the set of all possible solutions. The feasible solution space $\mathcal{S^{\prime}}\subseteq\mathcal{S}$ is the set of all feasible solutions, i.e., those satisfying problem constraints. The goal is to find the feasible solution $z^{*}\in\mathcal{S^{\prime}}$ that minimizes the cost function, i.e., $z^{*}=\arg\min_{z\in\mathcal{S^{\prime}}}C(z)$. A problem is said to be unconstrained if $\mathcal{S^{\prime}}=\mathcal{S}$, and constrained otherwise. The number of distinct cost values over the feasible solutions, i.e., the cardinality of the image of $C$ when restricted to the subdomain $\mathcal{S}^{\prime}$, is denoted by $m$. We denote the input size of a given problem instance as $s$  [^3] The input size is commonly denoted by $n$ in the literature, often coinciding with the number of vertices in a graph and/or the number of qubits. However, we adopt a different symbol in this work, as we deal with constrained problems where such a correspondence does not apply..
@@ 2.3 Quantum Walk Optimization Algorithm (QWOA)
  • $U_{W}(t_{l})=e^{-it_{l}H_{M}}$, where $H_{M}$ is the mixing Hamiltonian, performs a CTQW on a graph with an adjacency matrix $H_{M}$ over the feasible subspace $S^{\prime}$. Here, we set $H_{M}=K_{|\mathcal{S}^{\prime}|}$, where $K_{|\mathcal{S}^{\prime}|}$ represents the adjacency matrix of the complete graph.
@@
The $c$-approximate QWOA depth, for a constant $c$ and problem $\mathcal{P}$, denoted $p^{c}_{\mathcal{P}}$, is the minimum number of layers required by QWOA to achieves an approximation ratio $c$ for $\mathcal{P}$. We write $p^{c}$ to denote the corresponding quantity for any given problem.
@@ 2.4 Dynamical Lie Algebra (DLA)
Dynamical Lie algebras (DLAs) [22, 23, 24, 25, 26] are a useful tool for characterizing the expressiveness and trainability of PQCs. All operations on a quantum circuit acting on $n$-qubits correspond to an element of the special unitary group $\operatorname{SU}(2^{n})$, generated by the Lie algebra $\mathfrak{su}(2^{n})$. DLA has been shown to characterize the unitaries generated by parameterized quantum circuits of the form
$U(\boldsymbol{\theta})=\prod_{l=1}^{L}U_{l}(\boldsymbol{\theta}_{l}),\quad U_{l}(\boldsymbol{\theta}_{l})=\prod_{k=0}^{K}e^{-iH_{k}\theta_{lk}},$ (4)
@@
Definition 6 (Set of generators [22, Definition 2]).
Given a parametrized quantum circuit of the form in Eq. 4, we define the set of generators $\mathcal{G}=\{H_{k}\}_{k=0}^{K}$ as the set (of size $|\mathcal{G}|=K+1$) of the (traceless) Hermitian operators that generate the unitaries in a single layer of $U(\boldsymbol{\theta})$.
@@
Definition 7 (Dynamical Lie algebra (DLA) [22, Definition 3]).
@@
Let $\mathcal{G}$ be a set of generators (see Definition 6). The Dynamical Lie Algebra (DLA) $\mathfrak{g}$ is the subalgebra of $\mathfrak{su}(d)$ spanned by the repeated nested commutators of the elements in $\mathcal{G}$, i.e.,
$\mathfrak{g}=\langle iH_{o},\cdots,iH_{k}\rangle_{\operatorname{Lie}}\subseteq\mathfrak{su}(d),$ (5)
@@
Here, $\mathfrak{su}(d)$ is the special unitary algebra of degree $d$, i.e., the Lie algebra formed by the set of $d\times d$ skew-Hermitian, traceless matrices. $\langle S\rangle_{\operatorname{Lie}}$ denotes the Lie closure, i.e., the set obtained by repeatedly taking the nested commutators between the elements in $S$.
@@
We use $\mathfrak{g}_{\operatorname{QWOA},\mathcal{P}}$ to denote the DLA corresponding to the QWOA ansatz for an input problem $\mathcal{P}$. We write $\mathfrak{g}_{\operatorname{QWOA}}$ to indicate the DLA of the QWOA ansatz for any given problem. The Hamiltonians $H_{C}$ and $H_{M}$ form the set of generators for QWOA.
We say that $\mathfrak{g}_{j}\subseteq\mathfrak{g}$ is an ideal of $\mathfrak{g}$ if and only if $\mathfrak{g}_{j},\mathfrak{g}]\subseteq\mathfrak{g}_{j}$. The fact that $\mathfrak{g}$ is a reductive Lie algebra allows us to express $\mathfrak{g}$ as a direct sum of commuting ideals
$\mathfrak{g}=\mathfrak{g}_{1}\oplus\cdots\oplus\mathfrak{g}_{r},$ (6)
where $\mathfrak{g}_{j}\subseteq\mathfrak{g}$ is a simple Lie algebra (a Lie algebra that is non-abelian and contains no nontrivial ideals) for $j\in[r-1]$, $\mathfrak{g}_{r}$ is the center of $\mathfrak{g}$, and $\mathfrak{g},\mathfrak{g}]=\mathfrak{g}_{1}\oplus\cdots\oplus\mathfrak{g}_{r-1}$ is a semisimple, direct sum of simple Lie algebras. Readers are referred to Ref. [56] for a better understanding of Lie algebra.
@@
DLA allows one to explore desirable parameterized unitaries, addressing barren plateaus, overparameterization, and generalizationcapability [25, 23, 57]. From the lens of DLA, Ragone et al. [25] in their work identified conditions for which barren plateaus are present in a loss landscape (see more details in Section 2.5). Furthermore, Larocca et al. [23] showed that the number of parameters being larger than the dimension of the DLA is a sufficient condition for overparametrization for a PQC. Specifically, they showed that the dimension of the DLA is an upper bound on the threshold critical value for which a circuit is said to be overparameterized, and for the maximal rank attainable by the quantum Fisher information [58] and Hessian matrices [59]. In a recent work, Ohno [57] presented a generalization bound for quantum neural networks based on DLA. In particular, the generalization bound scales by $\mathcal{O}(\sqrt{\operatorname{dim}(\mathfrak{g})})$, where $\mathfrak{g}$ denotes the DLA of its generators.
2.5 Barren Plateaus
A major challenge for the trainability of VQAs is the phenomenon of barren plateaus (BPs) [21, 60, 61, 62], which arise in the broader context of PQCs. Barren plateaus correspond to regions on the optimization landscape where the variance of the loss function (and its gradients) vanishes exponentially with the system size. This flattens the optimization landscape, making gradient-based methods ineffective due to the lack of informative gradient signals. The BPs arise due to several factors such as expressiveness of the PQC [21], state initialization strategy [21], locality of the loss function measurement operator [61], and noise [60].
Given an initial state $\rho$, and an observable $O$, the loss function of a PQC $U(\boldsymbol{\theta})$ is given by
$L_{\boldsymbol{\theta}}(\rho,O)=\operatorname{Tr}[U(\boldsymbol{\theta})\rho U^{\dagger}(\boldsymbol{\theta})O],$ (7)
Ragone et al. [62] showed that the variance of the loss function is related to the DLA as follows
$\operatorname{Var}_{\boldsymbol{\theta}}[L_{\boldsymbol{\theta}}(\rho,O)]=\displaystyle\sum_{j=1}^{r-1}\frac{P_{j}(\rho)P_{j}(O)}{\operatorname{dim}(\mathfrak{g}_{j})},$ (8)
where $r$ is the total number of ideals into which the DLA can be decomposed. The exact value of $r$ depends on the particular DLA under consideration but is upper bounded by $\operatorname{dim}(\mathfrak{g})$. Here, the $\mathfrak{g}$-purity of a Hermitian operator $H\in i\mathfrak{u}(2^{n})$ is defined as [63, 64]
$P_{\mathfrak{g}}(H)=\operatorname{Tr}[H_{\mathfrak{g}}^{2}]=\sum_{j=1}^{\operatorname{dim}(\mathfrak{g})}|\operatorname{Tr}[E^{\dagger}_{j}H]|^{2},$ (9)
where $H_{\mathfrak{g}}$ is the orthogonal projection of $H$ onto $\mathfrak{g}_{\mathbb{C}}=\operatorname{span}_{\mathbb{C}}\mathfrak{g}$ (the complexification of $\mathfrak{g}$) and and $\{E_{j}:j\in[\operatorname{dim}(\mathfrak{g})]\}$ is an orthonormal basis for $\mathfrak{g}$. The loss is said to exhibit a BP if its variance vanishes exponentially with the system size, i.e., if $\operatorname{Var}_{\theta}[L_{\theta}(\rho,O)]\in\mathcal{O}(1/b^{n})$, for some $b>1$.
[…]
Note that QWOA deals with combinatorial optimization problems, where the most relevant parameter, the input size $s$ of the problem, does not necessarily coincide with the number of qubits — recall the constrained problems. In this context, we say that a QWOA loss function is said to exhibit barren plateaus if
[…]
for some $b>1$.

### v2

# Expressivity Limits in Quantum Walk-based Optimization

@@ Abstract
Quantum algorithms have emerged as a promising tool to solve combinatorial optimization problems. The quantum walk optimization algorithm (QWOA) is one such variational approach that has recently gained attention. In the broader context of variational quantum algorithms (VQAs), understanding the expressivity of the ansatz has proven critical for evaluating their performance. A key method to study this aspect involves analyzing the dimension of the dynamic Lie algebra (DLA). In this work, we derive novel upper bounds on the DLA dimension for QWOA applied to arbitrary optimization problems. Specifically, we show that the DLA dimension scales at most quadratically with the number of distinct eigenvalues of the problem Hamiltonian. As a consequence, our bound guarantees a polynomial DLA dimension with respect to the input size for optimization problems in the class $\mathsf{NPO}\text{-}\mathsf{PB}$. This result, coupled with recently established performance bounds for QWOA, allows us to identify complexity-theoretic conditions under which QWOA must be overparameterized to obtain optimal or approximate solutions for $\mathsf{NPO}\text{-}\mathsf{PB}$ problems.
@@ 1 Introduction
Combinatorial optimization problems play an important role in modeling and solving real-world systems. Some common classical approaches to solving these problems include Lagrange multipliers, simulated annealing, branch-and-bound, and evolutionary algorithms [1, 2]. Recent advances in quantum computing have also led to the development of quantum algorithms designed to solve these problems [3, 4, 5]. Among them, variational quantum algorithms (VQAs) [6] — which combine classical optimization with parametrized quantum circuits (PQCs), i.e., quantum circuits composed of unitary operations controlled by tunable parameters — have emerged as a prominent class and are particularly well suited for noisy quantum devices. Arguably, the most prominent VQA for combinatorial optimization is the quantum approximate optimization algorithm (QAOA) [7], a gate-based heuristic inspired by the quantum adiabatic algorithm [8, 9] that alternates between problem-specific and mixing unitaries to explore the solution space. A special variant of QAOA is the quantum walk optimization algorithm (QWOA) [10, 11], where the standard QAOA mixing unitary is replaced with a continuous-time quantum walk (CTQW) [12] operator. QWOA has been proposed as a natural alternative for solving combinatorial optimization problems with structured constraints, where quantum walks can be efficiently implemented through an indexing procedure that maps each feasible solution to a unique computational basis state. Typically, QWOA is implemented using a CTQW on a complete graph [13, 14, 15, 16], which is also the version we consider in this work  [^1] Note that, up to a rescaling of the variational parameter, QWOA is equivalent to a variant of QAOA called the Grover Mixer Quantum Alternating Operator Ansatz (GM-QAOA) [17]. Accordingly, all results and discussions in this work apply directly to that ansatz as well, and vice versa, meaning that results from that ansatz can be directly applied here..
@@
One of the key aspects for the success of a VQA is the expressivity of the PQC (a.k.a. ansatz), which should be sufficient to prepare the state that encodes the solution to the considered problem. Several approaches have been proposed to study the expressivity of PQCs, including techniques based on covering numbers [18, 19] and $t$-designs [20]. Yet another measure of expressivity, which we would focus on in this work, is the dimension of the dynamic Lie algebra (DLA) [21, 22, 23, 24, 25]. Given the set of Hamiltonians that generate the unitaries in a PQC, the corresponding DLA is constructed by taking repeated nested commutators of the Hamiltonians. This forms a Lie algebra that characterizes the space of unitaries accessible through time evolution. The dimension of the DLA determines the extent to which the circuit can explore the full unitary group or a restricted subspace, making it a fundamental measure of expressivity. If the DLA spans the full Lie algebra $\mathfrak{u}(d)$ for a $d$-dimensional Hilbert space, the ansatz is considered maximally expressive. However, in many practical scenarios, the DLA forms a lower-dimensional subalgebra, imposing structure on the reachable unitary transformations. As pointed out by Larocca et al. [22], the dimension of the DLA also provides an upper bound for the number of parameters required for a PQC to reach the overparameterization regime, beyond which additional parameters do not enhance the expressivity of the ansatz.
Beyond expressivity, the dimension of the DLA also influences other key aspects of VQAs, such as trainability and the classical simulability. In the vein of classical simulability, Goh et al. [26] showed that, given some additional conditions, if the DLA dimension is polynomial in the input size, one can efficiently compute expectation values of observables belonging to this algebra — as is the case for QAOA and QWOA. Regarding trainability, the dimension of the DLA is closely related to the emergence of barren plateaus [27] — regions in the parameter space where the variance of the loss function and its gradients vanishes exponentially with system size. In such regions, the optimization landscape becomes effectively flat, making gradient-based optimization methods inefficient and thus hindering the training of PQCs. In particular, Ragone et al. [24] established that the dimension of the DLA is inversely proportional to the variance of the loss function. This provides a rigorous framework for diagnosing the presence of barren plateaus based on algebraic properties of the circuit.
@@
In the context of QAOA, DLA has been studied both numerically and analytically [21, 22, 23, 25] for the Max-Cut problem, a combinatorial optimization problem defined on graphs and widely studied in the QAOA literature [28]. On the numerical side, the DLA dimension has been evaluated for two specific classes: $2$-regular graphs (also known as cycle graphs or ring of disagrees) and chain graphs (or path graphs), achieving scaling of $\mathcal{O}(n)$ and $\mathcal{O}(n^{2})$, respectively, where $n$ is the number of vertices of the graph [21, 23]. On the analytical side, Allcock et al. [25], using tools from group theory, proved that the exact dimension of the DLA for $2$-regular graphs is indeed linear in $n$, and also established that the exact DLA dimension for the complete graph is in the order of $n^{3}$. These findings, as discussed earlier, have implications in the context of expressivity of the QAOA ansatz for the Max-Cut problem. For instance, numerical findings by Larocca et al. [22] showed that the number of layers required to solve Max-Cut with a high success probability, on $2$-regular graphs and chains, matches the scaling of their corresponding DLA dimensions  [^2] Note that the number of variational parameters of QAOA/QWOA is linearly related to its number of layers. So, in asymptotic discussions, we can use the number of layers of QAOA/QWOA as a metric for parameterization without loss of generality., thus allowing us to infer that, for these classes of graphs, QAOA does not require overparameterization to achieve optimal solutions.
@@
While DLA has been extensively studied in the context of QAOA, it remains underexplored in the setting of QWOA. In this work, we initiate a systematic study of DLAs within the QWOA framework, focusing on whether overparameterization is necessary for solving or approximating combinatorial optimization problems.
@@
This paper is organized as follows. In Section 2, we introduce some background on combinatorial optimization, complexity classes, QWOA, and DLA. We present our main results and their theoretical implications in Section 3. Section 4 illustrates these results through the well-known problems of unstructured search, Max-Cut, and $k$-Densest Subgraph. Given the generality of the theoretical results presented, the examples included are not intended to serve as benchmarks, but rather to illustrate broader underlying phenomena. Lastly, we conclude our findings and discuss the future directions of our work in Section 5.
@@ 2.1 Combinatorial Optimization Problems
We consider combinatorial optimization problems defined by a cost function (the objective function) $C:\mathcal{S}\rightarrow\mathbb{R}$, where $\mathcal{S}$ — the solution space — is the set of all possible solutions. The feasible solution space $\mathcal{S^{\prime}}\subseteq\mathcal{S}$ is the set of all feasible solutions, i.e., those satisfying problem constraints. The goal is to find the feasible solution $z^{*}\in\mathcal{S^{\prime}}$ that minimizes the cost function, i.e., $z^{*}=\arg\min_{z\in\mathcal{S^{\prime}}}C(z)$. A problem is said to be unconstrained if $\mathcal{S^{\prime}}=\mathcal{S}$, and constrained otherwise. We denote the input size of a given problem instance as $s$  [^3] The input size is commonly denoted by $n$ in the literature, often coinciding with the number of vertices in a graph and/or the number of qubits. However, we adopt a different symbol in this work, as we deal with constrained problems where such a correspondence does not apply.. Let $\mathcal{C}$ be the image of $C$ when restricted to the subdomain $\mathcal{S}^{\prime}$. We denote the set $\mathcal{C}=\{x_{1},x_{2},\dots,x_{m}\}$, where each $x_{j}$ represents a distinct (feasible) cost value and $m=|\mathcal{C}|$. For $j=1,\ldots,m$, the respective multiplicity of each cost $x_{j}$ is denoted by $k_{j}$ (i.e., $k_{j}=|\{s\in\mathcal{S}^{\prime}\mid C(s)=x_{j}\}|$).
@@ 2.3 Quantum Walk Optimization Algorithm (QWOA)
  • $U_{W}(t_{l})=e^{-it_{l}H_{M}}$, where $H_{M}$ is the mixing Hamiltonian, performs a CTQW on a graph with an adjacency matrix $H_{M}$ over the feasible subspace $S^{\prime}$. Here, we consider a CTQW on a complete graph. Without loss of generality, we further take $H_{M}=J$, where $J$ is the all-ones matrix, since the resulting unitary operator is equivalent up to a global phase.
@@
The $c$-approximate QWOA depth, for a constant $c$ and problem $\mathcal{P}$, denoted $p^{c}_{\mathcal{P}}$, is the minimum number of layers required by QWOA to achieve an approximation ratio $c$ for $\mathcal{P}$. We write $p^{c}$ to denote the corresponding quantity for any given problem.
@@ 2.4 Dynamical Lie Algebra (DLA)
Dynamical Lie algebras (DLAs) [21, 22, 23, 24, 25] are a useful tool for characterizing the expressiveness of the PQCs. All operations on a quantum circuit acting on $d$-dimensional correspond to an element of the unitary group $\operatorname{U}(d)$, generated by the Lie algebra $\mathfrak{u}(d)$. DLA has been shown to characterize the unitaries generated by parameterized quantum circuits of the form
$U(\boldsymbol{\theta})=\prod_{l=1}^{L}U_{l}(\boldsymbol{\theta}_{l}),\qquad U_{l}(\boldsymbol{\theta}_{l})=\prod_{k=0}^{K}e^{-iH_{k}\theta_{lk}},$ (4)
@@
Definition 6 (Set of generators, adapted from Ref. [21], Definition 2).
Given a parametrized quantum circuit of the form in Eq. 4, we define the set of generators $\mathcal{G}=\{H_{k}\}_{k=0}^{K}$ as the set (of size $|\mathcal{G}|=K+1$) of the Hermitian operators that generate the unitaries in a single layer of $U(\boldsymbol{\theta})$  [^6] For simplicity, we do not impose the usual traceless condition on the generators, as this choice is convenient for the nature of our arguments. The only difference is the presence of identity terms, which merely introduce a global phase..
@@
Definition 7 (Dynamical Lie algebra (DLA), adapted from Ref. [21], Definition 3).
@@
Let $\mathcal{G}$ be a set of generators (see Definition 6). The Dynamical Lie Algebra (DLA) $\mathfrak{g}$ is the subalgebra of $\mathfrak{u}(d)$ spanned by the repeated nested commutators of the elements in $\mathcal{G}$, i.e.,
$\mathfrak{g}=\langle iH_{0},\cdots,iH_{K}\rangle_{\operatorname{Lie}}\subseteq\mathfrak{u}(d),$ (5)
@@
Here, $\mathfrak{u}(d)$ is the unitary Lie algebra of degree $d$, i.e., the Lie algebra formed by the set of $d\times d$ skew-Hermitian matrices. $\langle S\rangle_{\operatorname{Lie}}$ denotes the Lie closure, i.e., the set obtained by repeatedly taking the nested commutators between the elements in $S$.
@@
We use $\mathfrak{g}_{\operatorname{QWOA},\mathcal{P}}$ to denote the DLA corresponding to the QWOA ansatz for an input problem $\mathcal{P}$ and $\mathfrak{g}_{\operatorname{QWOA}}$ to indicate the DLA of the QWOA ansatz for any given problem. The Hamiltonians $H_{C}$ and $H_{M}$ form the set of generators for QWOA.
@@
Larocca et al. [22] showed that the number of parameters being larger than the dimension of the DLA is a sufficient condition for overparametrization for a PQC. Specifically, they showed that the dimension of the DLA is an upper bound on the threshold critical value for which a circuit is said to be overparameterized, and for the maximal rank attainable by the quantum Fisher information [57] and Hessian matrices [58]. Readers are referred to Ref. [59] for a better understanding of Lie algebra.